OpenStudy (anonymous):

Can someone show all the steps? Medal + fan My teacher wants me to show all steps and I dont even remember what I did. LOL thank you! a=b=c c^2 = a^2 + b^2 - 2ab cos(t) a^2 = a^2 + a^2 - 2a^2 cos(t) a^2 = 2a^2 - 2a^2 cos(t) 1 = 2[1 - cos(t)] cos(t) = 1/2 t=60

OpenStudy (anonymous):

do you know? @lol2214

OpenStudy (anonymous):

I'm trying to figure it out if I do, I'll post

OpenStudy (anonymous):

ok thank you so much (:

OpenStudy (anonymous):

Okay, so what exactly is it your teacher wants to know? Is this what you did above or is that the problem? Just the steps?

OpenStudy (anonymous):

My teacher asked to show how I got 60 as my answer and to show all my steps

OpenStudy (anonymous):

The law of cosines says c^2 = a^2 + b^2 - 2ab cos(t) In an equilateral triangle, a=b=c a^2 = a^2 + a^2 - 2a^2 cos(t) a^2 = 2a^2 - 2a^2 cos(t) 1 = 2[1 - cos(t)] cos(t) = 1/2 And t = 60 degrees when cos(t) = 1/2

OpenStudy (anonymous):

is that what i should tell her? @lol2214

OpenStudy (anonymous):

I'm pretty sure the link I sent you may help you decide from what your teacher is looking for.